MR-SDCI + Q AB INITIO MOLECULAR ORBITAL CALCULATIONS OF FeCO: IMPORTANCE OF WELL-CONTRIVED SA-MCSCF WAVE FUNCTIONS AND 8-10σ ELECTRON CORRELATIONS

AMANO MICHIKO, SACHIKO S. ITONO and \underline{TSUNEO HIRANO}, Department of Chemistry, Faculty of Science, Ochanomizu University, Tokyo 112-8610, Japan; UMPEI NAGASHIMA, Tsukuba Advanced Computing Center, National Institute of Advanced Industrial Science and Technology, Ibaraki 305-8562, Japan; MASAHIRO SEKIYA and KIYOSHI TANAKA, Division of Chemistry, Graduate School of Science, Hokkaido University, Sapporo 060-0810, Japan.

FeCO has been used as a benchmark molecule to evaluate the basis functions for Fe^{a}^{,}^{b} and the methods of calculation since it is known to be difficult to reproduce the energy difference between the ã^{5}Σ^{−} and the X̃^{3}Σ^{−} states as well as the experimentally observed bond lengths of the X̃^{3}Σ^{−} state by ab initio molecular orbital calculations. We carried out the MR-SDCI + Q and MR-ACPF calculations, based on the state-averaged MCSCF orbitals, taking into account the electron-correlation of 8-10σ electrons with the active space consisting of Fe 3d, 4s orbitals and CO π, π^{∗} orbitals. Our predicted term value of the ã^{5}Σ^{−} state, bond lengths r_{e}(Fe-C) and r_{e}(CO) of the X̃^{3}Σ^{−} state are 0.87 kcal mol^{−1}, 1.720 Å, and 1.159 Åwith relativistic energy corrections, which are to be compared with the corresponding experimental values of 3.24 kcal mol^{−1},^{c} 1.7270 Å [r_{s}(Fe-C)],^{d} and 1.1586 Å [r_{s}(CO)],^{d} respectively. Similar results have also been obtained by the MR-ACPF methods.

aA. Ricca and C. W. Bauschlicher, Theor Chem. Acc. \underline{106}, 314, (2001).

bT. Noro, M. Sekiya, T. Koga, and H. Matsuyama, Theor Chem. Acc. \underline{104}, 146, (2000).

cP. W. Villalta and D. G. Leopold, J. Chem. Phys. \underline{98}, 7730, (1993).

dK. Tanaka, S. Sakaguchi, and T. Tanaka, J. Chem. Phys. \underline{106}, 2118, (1997).